OpenStudy (anonymous):

What is the volume of the triangular prism? A) 480 cm3 B) 960 cm3 C) 1200 cm3 Eliminate D) 1920 cm3

OpenStudy (michele_laino):

hint: we can apply this formula: \[V = \frac{{Sh}}{3}\] where \(h=10\) cm, and \(S\) is the area of the base, which is the triangle, whose sides are \(12,\;16\) cm namely, I'm thinking this: 

OpenStudy (michele_laino):

oops.. I have ade an error, here is the right formula: \[V = Sh\]

OpenStudy (michele_laino):

so, we have: \[S = \frac{{12 \times 16}}{2} = ...?\]

OpenStudy (michele_laino):

please continue

OpenStudy (michele_laino):

oops.. I have made*...

OpenStudy (anonymous):

Um, so what is going on here I was in my phone sorry? All these errors???

OpenStudy (michele_laino):

as I wrote before, the right formula, is: \[V = Sh\] where: \[S = \frac{{12 \times 16}}{2} = 96\;c{m^2}\] and \(h=10\) cm therefore: \[V = Sh = 96 \times 10 = ...?\]

OpenStudy (michele_laino):

so, what is the right option?

OpenStudy (anonymous):

b